Circular No. 7054

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998et IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 18) on a red

plate taken on Nov. 17 UT with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. SN 1998et is located at α = 6^{h}59^{m}18^{s}.74, δ = +57^{o}44′41′′.0 (equinox 2000.0), which is 5′′.4 west and 5′′.4 north of the host galaxy’s center. There is no object at this position on a Palomar Sky Survey red plate taken on 1989 Dec. 5 or on the Digitized Sky Survey. S. Nakano, Sumoto, Japan, reports the following position end figures measured by Y. Kushida from an unfiltered CCD image taken in bad seeing by R. Kushida, Yatsugatake South Base Observatory, with a 0.40-m f/6.3 Schmidt-Cassegrain telescope on Nov. 18.773 UT, when the supernova’s magnitude was measured as 18.0: 18^{s}.36, 41′′.2. P. Garnavich, P. Challis, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 1998et was obtained by P. Berlind with the 1.5-m Tillinghast telescope on Nov. 19.4 UT, showing this to be a type-II supernova exhibiting strong Balmer and helium emission lines, but with a weak continuum and no P-Cyg absorption. The Hα line has a width of 1700 km/s (FWHM), and the supernova is a member of the peculiar type-IIn subclass. The host galaxy has a redshift of 12 100 km/s, based on the observed wavelength of narrow [O III] 500.7-nm emission.

SUPERNOVA 1998es IN NGC 632 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1998es was taken by M. Calkins on Nov. 14.2 UT at the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph). The spectrum features a blue continuum with strong absorption features at 374, 429, and 498 nm, and a weak, broad feature at 610 nm. The spectrum is generally similar to that of the peculiar type-Ia SN 1991T seven days before maximum light (Filippenko et al. 1992, Ap.J. 384, L15), with the exception of the absorption at 374 nm. The 374nm absoprtion was seen in the SN 1991T-like supernova 1998ab (IAUC 6858) and is presumably due to Ca II H and K. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 3168 km/s for NGC 632 implies a supernova expansion velocity of 19 000 km/s, measured using the Ca II H and K feature. The NED morphological classification of NGC 632 (Mrk 1002) is ‘S0?’, but it has a nuclear starburst with a diameter of 18′′ (Pogge and Eskridge 1993, A.J. 106, 1405), and the position of SN 1998es is projected near the edge of the star-forming region.”

1998 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green